OpenStudy (anonymous):

Area of regular polygons Find the area round to nearest 10th

OpenStudy (anonymous):

heres a picture http://gyazo.com/3b566c1f90063d81573556244431cb90

OpenStudy (anonymous):

nuber 18

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Regular triangle = equilateral triangle = all sides are the same length. Equilateral triangles can be split into two 30-60-90 degree right triangles. All you need is the long side of it, which will translate to the height of the original triangle. Because it's a 30-60-90 triangle, the ratios are from shortest to longest sides: x, x*sqrt(3), 2x. 2x=8, so x=4. The height ratio is x*sqrt(3), so h=4sqrt(3). Plug that into your triangle area formula, A=1/2bh. A=1/2(8)(4sqrt(3)) in^2 A=4(4sqrt(3)) in^2 A=16sqrt(3) in^2 Your final area is approximately 27.71 in^2, rounded.

OpenStudy (anonymous):

@tomhue i still dont get it on my homework its true or false and it asks true or false is number 18) 16

OpenStudy (anonymous):

Is the area 16 in^2? No, you'd mark that false.

OpenStudy (anonymous):

can you help me on my last one @tomhue

OpenStudy (anonymous):

http://gyazo.com/8324a4990944271f66035e981ccdff87 number 22